Microsoft's stock price peaked at 6118% of its I.P.O. price more than 13 years after the I.P.O.† Suppose that $45,000 invested in Microsoft at its I.P.O. price had been worth $2,700,000 (6000% of the I.P.O. price) after exactly 13 years. What interest rate, compounded annually, does this represent? (Round your answer to two decimal places.)
Answer by Boreal(15235)

2,700,000=45000(1+r)^13
60=(1+r)^13
logs of both sides
1.78=13 log (1+r)
round at end, not even here.
Divide by 13 then raise to the 10th power
1.370=1+r
subtract 1
0.37=r or 37%.
1.37^13=59.89, so checks with rounding.